OpenStudy (anonymous):

Find 3 positive values for "c" so that the trinomial is factorable 11y^2-40y+c

OpenStudy (whpalmer4):

11 is prime, right?

OpenStudy (anonymous):

yes..

OpenStudy (whpalmer4):

it is. that means that we're going to have something like \[(11x + a)(x - b)\]as our factoring.

OpenStudy (whpalmer4):

now, we need to find a pair of factors \(f_1, f_2\) such that \(11*f_1 + f_2=-40\) and \(f_1*f_2 = c\)

OpenStudy (whpalmer4):

one example would be -3 and -7. 11*-3 = -33, -33 + -7 = -40. -3*-7 = 21 \[(11y-7)(y-3) = 11y^2-33y-7y+21 = 11y^2-40y+21\]

OpenStudy (anonymous):

Thank you whpalmer4